# Are Data Augmentation Methods in Named Entity Recognition Applicable for Uncertainty Estimation?

### v1

@@ 1 Introduction
Our experiments yield several findings. First, some data augmentation methods in NER lead to improved confidence calibration and uncertainty estimation, especially in-domain. In particular, entity-prediction-based data augmentation Zhou et al. (2022) and entity replacement from the same entity type Dai and Adel (2020) show good performance. On the other hand, common confidence calibration methods, MC Dropout or TS Guo et al. (2017) have worse confidence calibration and uncertainty estimation performance than the data augmentation methods in NER, even though the data augmentation methods do not aim to improve confidence calibration and uncertainty estimation. Moreover, increasing the augmentation size improves performance in confidence calibration and uncertainty estimation. The improvement tends to be better the lower the perplexity of the sentences generated by the data augmentation. Our code will be released after acceptance.
@@ 5.2 Training Details
For the temperature parameter in TS, we used Optuna Akiba et al. (2019) to optimize the temperature parameter based on dev set loss with a search range of [0.001, 0.002, …, 5.000] in 100 trials. In addition, we optimized the binomial distribution parameter for data augmentation methods using the dev set by a grid search in the range of [0.1, 0.2, …, 0.8]. In LS, we conducted a grid search in the range of [0.01, 0.05, 0.1, 0.2, 0.3] to optimize the smoothing parameter. In the case of MELM, mask rate $\eta$ during fine tuning and mask parameter $\mu$ during generation are hyperparameters. We conducted a grid search for each hyperparameter in the range [0.3, 0.5, 0.7], as in Zhou et al. (2022).
@@
Table 2: Results of existing methods and data augmentation methods in OntoNotes 5.0 in ID setting. The best results are shown in bold. † indicates significantly improved than existing methods (p < 0.05).
@@ 6.1 Cross-genre Evaluation
AUPRC scores are shown in Table 4. Among the existing methods, TS shows superior performance; in data augmentation methods, MELM is not as good as in the case of calibration metrics such as ECE and MCE, and MR tends to show superior uncertainty performance.
@@ 6.2 Cross-lingual Evaluation
The results of cross-lingual transfer in MultiCoNER are shown in Table 5 with English as the source language. MR performs better in uncertainty performance for the ID situation. In contrast to the calibration and uncertainty performance in the cross-genre setting, both MR and SR show better calibration and uncertainty in the OOD setting. In Jiang et al. (2022), the result shows that the larger the linguistic distance Chiswick and Miller (2005), the more lenient the calibration and uncertainty estimation tends to be, and similar trends are obtained in this experiment. Unlike the discussion in Section 6.1, the uncertainty performance by data augmentation is also good for OOD in cross-lingual setting because the areas where only target set exist is limited in MultiCoNER (illustrated in Appendix F). On the other hand, MELM, which tends to show excellent calibration performance in cross-genre calibration, does not show good performance in cross-lingual settings.
@@ 6.3.1 Impact of Augmentation Size
To investigate the impact of data augmentation size on calibration and uncertainty performance, we analyze the trend of evaluation metrics in $\mathtt{tc}$$\mathtt{mz}$ scenario of OntoNotes 5.0 and $\mathtt{EN}$$\mathtt{ES}$ scenario of MultiCoNER, respectively. Figure 2 and 3 illustrate the results in the ID and OOD settings, respectively. In many cases, MR improves the calibration and uncertainty performance by increasing data. SR consistently improves as the dataset size doubles, whereas LwTR demonstrates only marginal improvement or even worsens as the dataset size increases. Finally, MELM improves further for OntoNotes 5.0 $\mathtt{tc}$, which shows excellent performance, and deteriorates further for MultiCoNER $\mathtt{EN}$, which shows poor performance.
@@ 6.3.2 Impact of Perplexities for Augmented Sentences
To investigate the influence of replacement units on data augmentation for NER as mentioned in Section 6.3.1, we measured the perplexity of the augmented sentences using GPT-2 Radford et al. (2019). The average perplexities of the augmented sentences and the average perplexities of the original training set for each dataset are shown in Table 6. Lower perplexity from augmented sentences tends to improve calibration performance and uncertainty performance. Consistently, the average perplexity of the sentences generated by MR is the lowest. Since MR performs substitutions on an entity-by-entity basis and does not affect the structure of the sentence itself, it has the lowest perplexity among the data augmentation methods in NER. MELM has the second lowest perplexity after MR, and may be adversely affected by generated entities that are adapted to the context but not actually present.
@@ 7 Conclusion
First, due to resource limitations, the experiment was limited to evaluation with English as the source language. To effectively investigate the calibration and uncertainty of zero-shot cross-lingual transfer, it is important to expand the investigation to include a wider range of languages as the source language. Therefore, future research should prioritize the investigation of calibration and uncertainty performance using different languages as the source for zero-shot cross-lingual transfer.
Second, as mentioned in Section 5.3, regarding the calibration and uncertainty evaluation policy, we simply evaluated an entity span as a single data instance, but a rigorous evaluation method that performs evaluation while considering multiple span candidates has been proposed Jiang et al. (2022). Establishing span-level NER calibration evaluation methods that can efficiently and comprehensively evaluate calibration and uncertainty for entity types for datasets with many entity types and long sequence lengths is a topic for future research.
Lastly, we broadly evaluated the calibration and uncertainty performance in both cross-genre and cross-lingual settings on data augmentation for NER, but only using sequence labeling-based methods. Recently, other paradigms in NER, such as span-based methods Fu et al. (2021) and Seq2Seq (sequence-to-sequence)-based methods Yan et al. (2021), have been proposed. In the future, the calibration or uncertainty performance of these methods could be evaluated.
@@ Acknowledgements
Appendix A Inference Time
@@
Methods OntoNotes 5.0 ($\mathtt{bc}$) OntoNotes 5.0 ($\mathtt{bn}$)
Methods OntoNotes 5.0 ($\mathtt{nw}$) OntoNotes 5.0 ($\mathtt{tc}$)
Table 8: F1 scores of existing calibration methods and data augmentation methods in OntoNotes 5.0.
Appendix B Details of Existing Calibration Methods
B.1 Temperature Scaling (TS)
$p_{i}=\frac{\exp(\bm{z}_{i}/T)}{\sum_{j=1}^{K}\exp(\bm{z}_{j}/T)}.$
B.2 Label Smoothing (LS)
$y^{LS}_{i}=y_{i}(1-\epsilon)+\frac{\epsilon}{K}.$
@@ B.3 Monte-Carlo Dropout (MC Dropout)
Methods $\mathtt{EN}$ $\mathtt{DE}$ $\mathtt{ES}$ $\mathtt{HI}$
Baseline 68.80$\pm$0.38 64.91$\pm$0.60 63.53$\pm$0.41 37.33$\pm$3.77
TS 68.51$\pm$0.52 64.70$\pm$0.90 63.41$\pm$0.45 37.90$\pm$2.79
LS 69.17$\pm$0.55 65.37$\pm$0.51 63.83$\pm$0.32 39.93$\pm$3.50
MC Dropout 68.56$\pm$0.96 64.70$\pm$0.87 63.39$\pm$0.69 36.38$\pm$5.89
LwTR (DA) 68.86$\pm$0.82 64.95$\pm$0.64 63.52$\pm$0.85 38.24$\pm$3.11
MR (DA) 69.71$\pm$0.72 65.37$\pm$0.57 64.25$\pm$0.62 37.53$\pm$4.03
SR (DA) 68.81$\pm$0.41 64.75$\pm$0.86 63.85$\pm$0.46 42.31$\pm$1.45
MELM (DA) 68.57$\pm$0.54 63.40$\pm$0.49 62.76$\pm$0.64 37.78$\pm$3.16
Table 9: F1 scores of existing calibration methods and data augmentation methods in MultiCoNER.
Appendix C F1 Scores
Table 8 and 9 show F1 scores. Note that in many cases, data augmentation methods do not degrade predictive performance itself, but MELM often significantly degrades predictive performance in some cases, especially when the source domains are $\mathtt{nw}$ and $\mathtt{tc}$. Considering Section 6.1 and 6.2, MR improves calibration and uncertainty performance in many cases without degrading predictive performance.
Appendix D More Results about Test Set Duplication
@@
Table 12: The percentage increase in entity duplication in the case of the source domain $\mathtt{tc}$ that are new overlaps with each target domain’s test set when applying each data augmentation method except MR. More results are in Appendix D.
Appendix E Impact of New Entities via Data Augmentation
@@
Appendix F t-SNE Plot for MultiCoNER Dataset
@@
Appendix G Results for Low-resource Language
@@
Appendix H Licenses of Datasets

### v2

@@ 1 Introduction
Our experiments yield several findings. First, some data augmentation methods in NER lead to improved confidence calibration and uncertainty estimation, especially in-domain. In particular, entity-prediction-based data augmentation Zhou et al. (2022) and entity replacement from the same entity type Dai and Adel (2020) show good performance. On the other hand, common confidence calibration methods, MC Dropout or TS Guo et al. (2017) have worse confidence calibration and uncertainty estimation performance than the data augmentation methods in NER, even though the data augmentation methods do not aim to improve confidence calibration and uncertainty estimation. Moreover, increasing the augmentation size improves performance in confidence calibration and uncertainty estimation. The improvement tends to be better the lower the perplexity of the sentences generated by the data augmentation. Our code is available on https://github.com/wataruhashimoto52/ner_da_uncertainty.
@@ 5.1 Datasets
Dataset &\&& Domain LwTR MR SR MELM ($\eta$, $\mu$)
OntoNotes 5.0
$\mathtt{bc}$ 0.3 0.7 0.3 (0.5, 0.5)
$\mathtt{bn}$ 0.4 0.8 0.2 (0.7, 0.3)
$\mathtt{mz}$ 0.7 0.4 0.5 (0.3, 0.3)
$\mathtt{nw}$ 0.7 0.5 0.7 (0.7, 0.7)
$\mathtt{tc}$ 0.4 0.4 0.1 (0.3, 0.3)
$\mathtt{wb}$ 0.7 0.7 0.8 (0.5, 0.7)
MultiCoNER
English ($\mathtt{EN}$) 0.2 0.8 0.4 (0.3, 0.3)
Table 2: Optimized hyperparameters in data augmentation methods in each source domain. We present the binomial distribution parameters for LwTR, SR and MR, and ($\eta$, $\mu$) for MELM, respectively.
@@ 5.2 Training Details
For the temperature parameter in TS, we used Optuna Akiba et al. (2019) to optimize the temperature parameter based on dev set loss with a search range of [0.001, 0.002, …, 5.000] in 100 trials. In addition, we optimized the binomial distribution parameter to manipulate replacement intensity for data augmentation methods using the dev set by a grid search in the range of [0.1, 0.2, …, 0.8]. In LS, we conducted a grid search in the range of [0.01, 0.05, 0.1, 0.2, 0.3] to optimize the smoothing parameter. In the case of MELM, mask rate $\eta$ during fine tuning and mask parameter $\mu$ during generation are hyperparameters. We conducted a grid search for each hyperparameter in the range [0.3, 0.5, 0.7], as in Zhou et al. (2022). All hyperparameters in data augmentation are shown in Table  2. The implementations of LwTR, MR and SR are used several repos, [^2] https://github.com/boschresearch/data-augmentation-coling2020  [^3] https://github.com/kajyuuen/daaja while the implementation of MELM used the official repo. [^4] https://github.com/RandyZhouRan/MELM
@@
Table 3: Results of existing methods and data augmentation methods in OntoNotes 5.0 in ID setting. The best results are shown in bold. † indicates significantly improved than existing methods (p < 0.05) by using t-test.
@@ 6.1 Cross-genre Evaluation
AUPRC scores are shown in Table 5. In the AUPRC scores in OntoNotes 5.0, data augmentation methods are outperform existing methods in 15 cases out of 24 cases. Among the existing methods, TS shows superior performance; in data augmentation methods, MELM is not as good as in the case of calibration metrics such as ECE and MCE, and MR tends to show superior uncertainty performance. Calibration and scores based on AUC measure different points of uncertainty Galil et al. (2023), therefore we assume that uncertainties that can be improved vary depending on the methods.
@@ 6.2 Cross-lingual Evaluation
The results of cross-lingual transfer in MultiCoNER are shown in Table 6 with English as the source language. MR performs better in uncertainty performance for the ID situation. In contrast to the calibration and uncertainty performance in the cross-genre setting, both MR and SR show better calibration and uncertainty in the OOD setting. In Jiang et al. (2022), the result shows that the larger the linguistic distance Chiswick and Miller (2005), the more lenient the calibration and uncertainty estimation tends to be, and similar trends are obtained in this experiment. Unlike the discussion in Section 6.1, the uncertainty performance by data augmentation is also good for OOD in cross-lingual setting because the areas where only target set exist is limited in MultiCoNER (illustrated in Appendix G). On the other hand, MELM, which tends to show excellent calibration performance in cross-genre calibration, does not show good performance in cross-lingual settings.
@@ 6.3.1 Impact of Augmentation Size
To investigate the impact of data augmentation size on calibration and uncertainty performance, we analyze the trend of evaluation metrics in $\mathtt{tc}$$\mathtt{mz}$ scenario of OntoNotes 5.0 and $\mathtt{EN}$$\mathtt{ES}$ scenario of MultiCoNER, respectively. Figure 2 and 3 illustrate the results in the ID and OOD settings, respectively. In many cases, MR improves the calibration and uncertainty performance by increasing data. [^5] Note that we have not discussed about the absolute values of the uncertainty estimation performance. SR consistently improves as the dataset size doubles, whereas LwTR demonstrates only marginal improvement or even worsens as the dataset size increases. Finally, MELM improves further for OntoNotes 5.0 $\mathtt{tc}$, which shows excellent performance, and deteriorates further for MultiCoNER $\mathtt{EN}$, which shows poor performance.
@@ 6.3.2 Impact of Perplexities for Augmented Sentences
To investigate the influence of replacement units on data augmentation for NER as mentioned in Section 6.3.1, we measured the perplexity of the augmented sentences using GPT-2 Radford et al. (2019). The average perplexities of the augmented sentences and the average perplexities of the original training set for each dataset are shown in Table 7. Lower perplexity from augmented sentences tends to improve calibration performance and uncertainty performance. Consistently, the average perplexity of the sentences generated by MR is the lowest. Since MR performs substitutions on an entity-by-entity basis and does not affect the structure of the sentence itself, it has the lowest perplexity among the data augmentation methods in NER. [^6] As shown in Appendix I, not only the uncertainty performance but also the prediction performance could be affected by preserving the structure of a sentence. MELM has the second lowest perplexity after MR, and may be adversely affected by generated entities that are adapted to the context but not actually present.
@@ 7 Conclusion
Source Language
Due to resource limitations, the experiment was limited to evaluation with English as the source language. To effectively investigate the calibration and uncertainty of zero-shot cross-lingual transfer, it is important to expand the investigation to include a wider range of languages as the source language. Therefore, future research should prioritize the investigation of calibration and uncertainty performance using different languages as the source for zero-shot cross-lingual transfer.
Evaluation of Uncertainty for Entities
As mentioned in Section 5.3, regarding the calibration and uncertainty evaluation policy, we simply evaluated an entity span as a single data instance, but a rigorous evaluation method that performs evaluation while considering multiple span candidates has been proposed Jiang et al. (2022). Establishing span-level NER calibration evaluation methods that can efficiently and comprehensively evaluate calibration and uncertainty for entity types for datasets with many entity types and long sequence lengths is a topic for future research.
NER Paradigm
We broadly evaluated the calibration and uncertainty performance in both cross-genre and cross-lingual settings on data augmentation for NER, but only using sequence labeling-based methods. Recently, other paradigms in NER have been proposed such as the span-based methods Fu et al. (2021) and the generation-based methods Yan et al. (2021) including BART Lewis et al. (2020) or Large Language Models (LLM)  Xu et al. (2024), which are also applicable to nested-NER. In the future, the calibration or uncertainty performance of these methods could be evaluated.
Other Data Augmentation Methods
In this study, we focused on the data augmentation methods based on token or entity replacement. On the other hand, paraphrase-based data augmentation methods using such as LLM have attracted attention Ding et al. (2024). By using LLM, it is also possible to generate entities that correspond to a specified entity type Ye et al. (2024). To Investigate these in the context of uncertainty estimation also will be an interesting research.
@@ Acknowledgements
Appendix A Licenses of Datasets
Appendix B Details of Existing Calibration Methods
B.1 Temperature Scaling (TS)
$p_{i}=\frac{\exp(\bm{z}_{i}/T)}{\sum_{j=1}^{K}\exp(\bm{z}_{j}/T)}.$
B.2 Label Smoothing (LS)
$y^{LS}_{i}=y_{i}(1-\epsilon)+\frac{\epsilon}{K}.$
@@ B.3 Monte-Carlo Dropout (MC Dropout)
Methods $\mathtt{bc}$ $\mathtt{bn}$ $\mathtt{nw}$ $\mathtt{tc}$
ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓) ECE (↓) MCE (↓)
Baseline 27.07 33.52 26.08 31.17 26.66 31.35 37.66 46.32
TS 27.25 33.41 26.17 31.17 26.68 31.34 36.66 45.52
LS 27.19 33.57 25.88 30.49 26.52 30.67 35.24 43.68
MC Dropout 27.15 33.61 25.90 30.85 26.62 31.18 36.80 45.71
LwTR (DA) 27.65 33.78 26.49 31.78 27.28 31.67 35.90 44.97
MR (DA) 27.33 33.22 26.21 31.00 26.26 30.53 36.38 44.65
SR (DA) 27.23 33.08 26.11 30.72 27.47 31.89 35.24 43.57
MELM (DA) 27.95 33.88 26.63 30.91 27.62 32.09 34.83 42.65
Table 9: ECE and MCE averaged over all target domain results in OntoNotes 5.0.
Appendix C Inference Time
Table 8 shows the results of the inference time on MultiCoNER $\mathtt{EN}$ set. We can see that data augmentation methods do not affect the computational overhead during inference clearly.
Appendix D Full Averaged Results on OntoNotes 5.0
To briefly summarize the many values in Table 3 and  4, we averaged the ECE and MCE scores for each method and domain, shown them in Table 9. From this table, we can see that data augmentation methods are slightly worse than existing methods some cases when averaging all settings , while in others, especially $\mathtt{nw}$ and $\mathtt{tc}$, data augmentation methods are better on average.
Appendix E More Results about Test Set Duplication
@@
Table 12: The percentage increase in entity duplication in the case of the source domain $\mathtt{tc}$ that are new overlaps with each target domain’s test set when applying each data augmentation method except MR. More results are in Appendix E.
Appendix F Impact of New Entities via Data Augmentation
@@
Appendix G t-SNE Plot for MultiCoNER Dataset
@@
Appendix H Results for Low-resource Language
@@
Methods OntoNotes 5.0 ($\mathtt{bc}$) OntoNotes 5.0 ($\mathtt{bn}$)
Methods OntoNotes 5.0 ($\mathtt{nw}$) OntoNotes 5.0 ($\mathtt{tc}$)
Table 14: $F_{1}$ scores of existing calibration methods and data augmentation methods in OntoNotes 5.0.
Methods $\mathtt{EN}$ $\mathtt{DE}$ $\mathtt{ES}$ $\mathtt{HI}$
Baseline 68.80$\pm$0.38 64.91$\pm$0.60 63.53$\pm$0.41 37.33$\pm$3.77
TS 68.51$\pm$0.52 64.70$\pm$0.90 63.41$\pm$0.45 37.90$\pm$2.79
LS 69.17$\pm$0.55 65.37$\pm$0.51 63.83$\pm$0.32 39.93$\pm$3.50
MC Dropout 68.56$\pm$0.96 64.70$\pm$0.87 63.39$\pm$0.69 36.38$\pm$5.89
LwTR (DA) 68.86$\pm$0.82 64.95$\pm$0.64 63.52$\pm$0.85 38.24$\pm$3.11
MR (DA) 69.71$\pm$0.72 65.37$\pm$0.57 64.25$\pm$0.62 37.53$\pm$4.03
SR (DA) 68.81$\pm$0.41 64.75$\pm$0.86 63.85$\pm$0.46 42.31$\pm$1.45
MELM (DA) 68.57$\pm$0.54 63.40$\pm$0.49 62.76$\pm$0.64 37.78$\pm$3.16
Table 15: $F_{1}$ scores of existing calibration methods and data augmentation methods in MultiCoNER.
Appendix I $F_{1}$ Scores
Table 14 and 15 show $F_{1}$ scores. Note that in many cases, data augmentation methods do not degrade predictive performance itself, but MELM often significantly degrades predictive performance in some cases, especially when the source domains are $\mathtt{nw}$ and $\mathtt{tc}$. Considering Section 6.1 and 6.2, MR improves calibration and uncertainty performance in many cases without degrading predictive performance.